OpenStudy (anonymous):

Select the fraction that is equivalent to 42.8181

OpenStudy (helder_edwin):

it is \[ \large 42.\overline{81}=42+0.\overline{81} \]

OpenStudy (anonymous):

42.81

OpenStudy (helder_edwin):

if \(x=0.\overline{81}\) then \[ \large 100x=81.\overline{81} \] whence \[ \large 99x=100x-x=81 \] then \[ \large x=\frac{81}{99}=\frac{9}{11} \]

OpenStudy (anonymous):

Select one of the options below as your answer: A. 471 11 B. 42 81 C. 4,281 100 D. 42 81 100

OpenStudy (helder_edwin):

it is not 42.8181818181818181.....

OpenStudy (anonymous):

ik?

OpenStudy (helder_edwin):

well \[ \large 42.81=42+0.81=42+\frac{81}{100}=\frac{4281}{100} \]

OpenStudy (anonymous):

ok thanks

OpenStudy (helder_edwin):

u r w

OpenStudy (anonymous):

:) im a fan

OpenStudy (helder_edwin):

thanks